# Cherry on Top: Parameter Heterogeneity and Quantization in Large Language Models

### v1

@@ Abstract
This paper reveals the phenomenon of parameter heterogeneity in large language models (LLMs). We find that a small subset of ”cherry” parameters exhibit a disproportionately large influence on model performance, while the vast majority of parameters have minimal impact. This heterogeneity is found to be prevalent across different model families, scales, and types. Motivated by this observation, we propose CherryQ, a novel quantization method that unifies the optimization of mixed-precision parameters. CherryQ identifies and preserves the critical cherry parameters in high precision while aggressively quantizing the remaining parameters to low precision. Extensive experiments demonstrate the effectiveness of CherryQ. CherryQ outperforms existing quantization approaches in terms of perplexity and downstream task performance. Notably, our 3-bit quantized Vicuna-1.5 exhibits competitive performance compared to their 16-bit counterparts. These findings highlight the potential of CherryQ for enabling efficient deployment of LLMs by taking advantage of parameter heterogeneity.
@@ 1 Introduction
Quantizing parameters from higher bits to the integer points in the lower-bit space inevitably perturbs the parameters from its optimum, leading to a degradation in performance (i.e. quantization error). To mitigate this error, various approaches have been proposed, such as iterative block-wise quantization ([15, 10, 8]), gradient approximation ([18, 3]), and low-rank approximation of perturbations ([9]). However, existing approaches still cause clear performance degradation, especially for extreme low bits (e.g. 3-bit).
Before investigating how to further mitigate the performance degradation, we raise a more fundamental question: To what extent can quantization errors be mitigated? Our study shows that the answer is more complex than expected. For the vast majority ($>99\%$) of parameters, their quantization errors are minimal and thus can be alleviated or ignored. Nonetheless, there exists a small subset of parameters ($<1\%$) for which the quantization errors are substantial and hard to mitigate.
@@
Figure 1: Scatter plot of parameter impacts in different LLMs. We randomly sampled 4096 parameters from the corresponding parameter matrix. Each point represents the impact of an individual parameter. Insets show the zoomed-in y-axis. The heterogeneity is found in different model scales (1(a),1(b)), different model families (1(c), 1(d)), and both base models and chat models (1(e), 1(f)).
Consider Figure 1(a) as an example. We show the scatter plot of impacts on quantization error when perturbing each individual parameters in a parameter matrix from LLaMA2-7b [23]. The derivation of impacts are detailed in § 3. As 99% of parameters are in the range of (0,0.1), a small subset of “cherry” parameters exhibit a disproportionately large influence in the range of (5,30), with 50-300 times greater than the maximum value of the remaining 99% parameters.
This phenomenon is not an isolated occurrence. We observed similar patterns across different LLM scales (Figure 1(a)1(b)), different LLM families, including Mistral [12] (Figure 1(c)) and Gemma [22] (Figure 1(d)), and both base models and chat models (Vicuna-1.5 [5] Figure 1(e)1(f)). The consistent presence suggests that it is an inherent characteristic of LLMs. We will elaborate the evidences from a more macro view in § 5. Based on these findings, we introduce the phenomenon of parameter heterogeneity:
Parameter Heterogeneity in LLMs:
Consider the impact of parameter perturbations on the model’s behavior, a small set of “cherry” parameters have a disproportionately large influence. In contrast, the vast majority of normal parameters have a minimal impact.
We emphasize that the parameter impact is heterogeneous, as it exhibits a more severe imbalance compared to the commonly observed imbalance in LLMs [18, 17, 6, 7]. As a contrast, in § 5, we demonstrate that the parameter impact is much more heterogeneous than the parameter magnitude, although the latter one is also found to be imbalanced [7, 18]. This stark difference highlights the significance of considering parameter impact when optimizing and quantizing LLMs.
The parameter heterogeneity poses new challenges for conventional quantization strategies. Although quantizing most normal parameters will cause little performance degradation, cherry parameters are highly sensitive to quantization-induced perturbations. However, existing strategies usually fail to preserve the delicate structure of these critical parameters [8, 18].
Addressing this challenge is not trivial. One straightforward way is to employ a mixed-precision quantization [13], representing cherry parameters and normal parameters with high and low precisions, respectively. However, the simultaneous optimization of both types of parameters becomes a major challenge. In the widely used GPTQ [8] approach and PTQ framework, the optimal values of the early quantized parameters may change as the cherry parameters are updated. However, parameters cannot be updated once they are quantized in PTQ, which limits the early quantized parameters from reaching their optimal values.
To address this challenge, we employ the quantization-aware training (QAT) framework to handle parameter heterogeneity. Our quantization optimize the mixed-precision parameters with a unified backpropagation. For the cherry parameters, we maintain their 16-bit representations and apply standard gradient descent. For the normal parameters, we apply an extra Straight-Through Estimator (STE) trick [3] for gradient descent. Therefore, all parameters are continuously and uniformly optimized. We denote the approach as CherryQ (Cherry parameter and Quantization-aware training).
@@
Extensive experiments on different models and benchmarks demonstrate the efficacy of CherryQ. It consistently yields the lowest perplexity on most settings. Notably, our 3-bit Vicuna-1.5 model exhibit performance on par with the 16-bit counterpart on Vicuna-bench [5].
We believe that our work opens up new avenues for understanding and harnessing the complex interplay of parameters in LLMs. The heterogeneity offers a novel perspective to navigate the trade-off between parameter efficiency and performance.
@@ 2 Related Work
Quantization Strategies for LLMs Various quantization strategies have been proposed in the literature to reduce the precision of weights and activations while maintaining acceptable accuracy. These strategies can be broadly categorized into post-training quantization and quantization-aware training [14]. Post-training quantization methods, such as OBD, OBS, and GPTQ, directly quantize the pre-trained model without fine-tuning [15, 10, 8]. On the other hand, quantization-aware training methods, such as LLM-QAT [18], incorporate quantization operations into the training process to jointly optimize the quantized model. Some works also explore mixed-precision quantization [13] and adaptive quantization bins [7] to achieve a better trade-off between accuracy and efficiency.
Outliers in Language Model Quantization The idea of modeling parameter outliers in LLM quantization is not new. The exploration of outliers primarily includes the perspectives of magnitude [18, 7] and activations [4, 6]. For example, from the magnitude perspective, QLoRA assumes that parameters follow a Gaussian distribution [7] and designs information-theoretically optimal quantized bins based on this assumption. [18] keeps outlier parameters in 16-bit precision. From the activation perspective, [17] migrates the outlier amplifier to subsequent modules through an equivalent transformation. Additionally, SqueezeLLM also measures outliers from the perspective of parameter impact [13]. To the best of our knowledge, our work is the first to systematically reveal the outliers (heterogeneity) of parameter impact across different models, and we show a more pronounced imbalance in parameter impacts compared to magnitudes (§ 6.4). Furthermore, we propose a method to unify outlier (cherry) parameter optimization and normal parameter optimization, addressing the optimization challenges of heterogeneous parameters.
@@ 3 Quantifying the Impact of Parameters on Model Performance
$\mathbf{L}(w_{i}+\Delta)-\mathbf{L}(w_{i})=g_{i}\Delta+\frac{1}{2}\mathbf{H}_{ii}\Delta^{2}+O(|\Delta|^{3})$ (1)
@@
4 Unified Mixed-Precision Training
The insights gained from Figure 1 highlights the heterogeneity in model parameters. The cherry parameters, despite constituting less than 1% of the total parameter count, exert a substantial influence on the model. Indiscriminately quantizing these cherry parameters alongside the normal parameters may lead to a significant deterioration in model performance.
To mitigate the impact of cherry parameters on quantization, we propose to preserve their high-precision values during the quantization process. By maintaining the fidelity of these critical parameters, we ensure that the essential information they capture is not compromised.
Optimizing mixed-precision parameters in LLMs presents a unique challenge. The widely adopted GPTQ approach [8], which falls under the Post-Training Quantization (PTQ) framework [14], struggles to simultaneously optimize high-precision cherry parameters and low-precision normal parameters. This is because updating the cherry parameters during the PTQ process significantly affects the model, causing the optimal values of the normal parameters to vary. However, in the PTQ framework, once the parameters are quantized, they cannot be updated further. This limitation prevents the early-stage quantized parameters from reaching their optimal values. On the other hand, if we do not allow the updates of the cherry parameters during the PTQ process [17], the quantized model will lose the flexibility provided by these critical parameters.
To address this challenge, we propose a novel approach that unifies the optimization of mixed-precision parameters. Our method leverages a QAT framework, which allows for the simultaneous optimization of both cherry parameters and normal parameters. During backpropagation, the high-precision cherry parameters are updated using standard gradient descent, while the low-precision normal parameters employ the Straight-Through Estimator (STE) trick [3] for low precision gradient descent. This unified backpropagation enables end-to-end optimization of both cherry parameters and normal parameters, enhancing the overall optimization effect. We show the quantization in Algorithm 1.
Algorithm 1 Cherry Parameter and Quantization-aware Training (CherryQ)
@@ 4 Unified Mixed-Precision Training
3:$\mathbf{C}\leftarrow\{w_{i}\in\mathbf{W}\mid H_{ii}>\tau\}$triangleright Identify cherry parameters
@@
6:     $L\leftarrow\mathrm{model}(x;\mathbf{C}\cup Quant(\mathbf{N}))$triangleright Compute loss
7:     $\mathbf{C}\leftarrow\mathbf{C}-\eta\frac{\partial L}{\partial\mathbf{C}}$triangleright Standard gradient descent
8:     $\mathbf{N}\leftarrow\mathbf{N}-\eta\cdot\mathrm{STE}(\frac{\partial L}{\partial\mathbf{N}})$triangleright Gradient descent with gradient approximation by STE
@@
5 Prevalence of Parameter Heterogeneity in LLMs
While Figure 1 showcases the heterogeneity of selected parameter matrices in different LLMs, it is crucial to investigate whether this phenomenon is pervasive across the hundreds of parameter matrices within each LLM. In this section, we conduct a comprehensive analysis of parameter heterogeneity from a macro perspective.
To quantify the degree of heterogeneity in a parameter matrix, we introduce the heterogeneity score of the matrix. Inspired by the observation in Figure 1, where a small subset of parameters exhibits significantly higher impacts compared to the maximum of the majority, we define the heterogeneity score as the ratio of the mean impact of the top 1% parameters to the maximum impact of the bottom 99% parameters, as shown in Equation (4). A higher heterogeneity score indicates a more pronounced disparity in parameter importance within the matrix.
Heterogeneity Score=Mean($\text{Heterogeneity Score}=\frac{\text{Mean}(\mathbf{H}_{ii\text{ top 1\%}})}{\text{Max}(\mathbf{H}_{ii\text{ bottom 99\%}})}$ (4)
For comparison, we also include the heterogeneity scores based on the magnitude of parameters, a commonly used measure of parameter importance [11]. The magnitude-based heterogeneity score is calculated using Equation (5).
Magnitude Heterogeneity $\text{Magnitude Heterogeneity Score}=\frac{\text{Mean}(|w_{i}|_{\text{ top 1\%}})}{\text{Max}(|w_{i}|_{\text{ bottom 99\%}})}$ (5)
To provide a comprehensive view of parameter heterogeneity across different matrices, we plot the scatter distribution of heterogeneity scores for all parameter matrices of each model in Figure 2. It clearly reveals that the parameter matrices across different LLMs exhibit high heterogeneity scores, especially when comparing with parameter magnitudes. This finding strongly suggests that parameter heterogeneity is not an isolated occurrence but rather a widespread phenomenon in LLMs.
The pervasiveness of parameter heterogeneity highlights the need for quantization strategies that can effectively handle the disparate importance of parameters, ensuring that the cherry parameters are preserved with higher precision while allowing for more aggressive quantization of the less influential normal parameters.
@@ 5 Prevalence of Parameter Heterogeneity in LLMs
Figure 2: Scatter distribution of heterogeneity scores for different parameter matrices in LLMs. Each point represents a parameter matrix, with the x-axis indicates the matrix index and the y-axis showing the heterogeneity score.
@@ 6 Quantization Experiments
In the experimental section, we demonstrate the effectiveness of the CherryQ for both base LLMs and chat LLMs. We also compare different cherry parameter selection criteria to highlight the impact-based heterogeneity.
@@ 6.1 Implementation Details
Parameter Representation: Based on the observation that cherry parameters occupy a very small proportion, for each row of parameters in each parameter matrix, we only consider the top 1/256 parameters with the highest impact as cherry parameters and retain their FP16 precisions. For example, the parameter matrix size of LLaMA2-7B is $4096\times 4096$. So we select $16$ parameters with the highest impact for each row, which results in $4096\times 16$ parameters as cherry parameters. Additionally, to recover the complete parameter matrix, an INT16 is required to record the column index for each cherry parameter. So each cherry parameter need 32 bits.
[…]
$Dequant(S,X^{Intk})=S(X^{Intk}+0.5)$ (7)
Quantization Datasets: For the quantization of the base LLMs, we follow [8] to use C4 [20] as the training data. We select the first 4 partitions of C4 and choose data with a length of $\geq 2048$ tokens, resulting in a total of 50k samples of 2048 token. For the chat LLMs, since Vicuna-1.5 [5] is obtained by supervised fine-tuning based on ShareGPT [5], we also use the ShareGPT dataset for training. We utilize a total of 20k training samples from ShareGPT for QAT and Cherry.
Baselines We compare our method with various quantization methods, including QAT [18], GPTQ [8], SqueezeLLM [13], OminiQuant [21], and AWQ [17]. For OminiQuant and AWQ, we use their results reported in [21]. For SqueezeLLM, we use the results in its original paper [13]. For GPTQ, its 4-bit model is obtained from the open-source  [^1] https://huggingface.co/TheBloke. Due to the lack of a 3-bit GPTQ model, we quantize the model ourselves via the implementation of Auto-GPTQ  [^2] https://github.com/AutoGPTQ/AutoGPTQ. Since CherryQ is based on QAT, for fair comparisons, the implementation of QAT is the same as CherryQ, except that it does not handle cherry parameters.

### v2

@@ Abstract
This paper reveals the phenomenon of parameter heterogeneity in large language models (LLMs). We find that a small subset of “cherry” parameters exhibit a disproportionately large influence on model performance, while the vast majority of parameters have minimal impact. This heterogeneity is found to be prevalent across different model families, scales, and types. Motivated by this observation, we propose CherryQ, a novel quantization method that unifies the optimization of mixed-precision parameters. CherryQ identifies and preserves the critical cherry parameters in high precision while aggressively quantizing the remaining parameters to low precision. Extensive experiments demonstrate the effectiveness of CherryQ. CherryQ outperforms existing quantization approaches in terms of perplexity and downstream task performance. Notably, our 3-bit quantized Vicuna-1.5 exhibits competitive performance compared to their 16-bit counterparts.
@@ 1 Introduction
Figure 1: Scatter plot of parameter impacts in different LLMs. We randomly sampled 4096 parameters from the corresponding parameter matrix. Each point represents the impact of an individual parameter. Insets show the zoomed-in y-axis. The heterogeneity is found across different model scales (1(a),1(b)), different model families (1(c), 1(d)), and both base models and chat models (1(e), 1(f)).
Low-precision parameter representation leads to quantization errors. Surprisingly, existing research has shown that LLMs exhibit a high robustness for quantization errors even for low-bit settings. For example, although 4-bit quantization can only represent 16 distinct values, even the simplest round-to-nearest strategy does not significantly degrade performance [17]. This raises the question: what causes LLMs to be robust to quantization?
We explore the parameters and answer this question via Parameter Heterogeneity, which refers to the significant variation in the influence of quantization on different parameters. We reveal that for the vast majority ($>99\%$) of normal parameters, the effect of their quantization to the model are minimal and can thus be alleviated or ignored. However, there exists a small subset ($<1\%$) of “cherry” parameters for which the effect are substantial and hard to mitigate.
Consider Figure 1(a) as an example. We show a scatter plot of the impacts on the model loss when perturbing each individual parameter in a parameter matrix from LLaMA2-7b [23]. The derivation of impacts is detailed in § 3. While 99% of the parameters fall within the range of (0, 0.1), a small subset of “cherry” parameters exhibits values ranging from (5, 30), which is 50-300 times higher than the maximum value of the remaining 99% of parameters. Moreover, this phenomenon is not an isolated case. We observed similar patterns across different scales of LLMs (Figure 1(a)1(b)), different families of LLMs, including Mistral [11] (Figure 1(c)) and Gemma [22] (Figure 1(d)), and both base models and chat models (Vicuna-1.5 [5] Figure 1(e)1(f)). The consistent presence of this phenomenon suggests that parameter heterogeneity is an inherent characteristic of LLMs.
Therefore, 99% of normal parameters explain the high robustness of LLMs to quantization errors. However, the small number of cherry parameters still leads to performance degradation under quantization. Consequently, the key to reducing quantization errors lies in addressing the quantization of cherry parameters.
The parameter heterogeneity also explains the previously discovered effectiveness of mixed-precision quantization strategies [6, 12, 17, 18]. By preserving a small proportion of parameters with high precision, the quantization performance can be effectively improved. Based on the heterogeneity, this strategy alleviates the impact of cherry parameters on model performance by maintaining their precision.
Indeed, mixed-precision strategies can effectively address the quantization error issue of cherry parameters. However, the core challenge lies in identifying cherry parameters based on specific metrics. Different metrics of parameter effects include weights [6, 13, 18], activations [17, 24], output changes [8], as well as the impact of parameters on model loss used in this paper. Based on the parameter heterogeneity, we argue that effective cherry parameter identification metrics should exhibit high heterogeneity, clearly distinguishing between cherry parameters and normal parameters. Accordingly, we compare three different metrics in Sec 5 and find that the impact best distinguishes cherry parameters from normal parameters. The experimental results in Sec 6.5 verify that choosing metrics with higher discriminative capability indeed leads to better performance.
@@
Based on the above analysis, we design the CherryQ quantization algorithm, which selects cherry parameters based on the impact metric and end-to-end optimizes parameters with mixed precisions. Extensive experiments on various models and benchmarks demonstrate the effectiveness of CherryQ. It consistently yields the lowest perplexity in most settings. Notably, our 3-bit Vicuna-1.5 model exhibits performance on par with the 16-bit counterpart on Vicuna-bench [5]. Our 2-bit quantization method significantly outperforms the SOTA approaches.
In summary, by systematically revealing parameter heterogeneity in LLMs, we answer the following questions:
  1. 1.
    What causes the high robustness of LLMs to quantization? It is due to the 99% of normal parameters in parameter heterogeneity.
  2. 2.
    Why is mixed-precision quantization effective? This strategy addresses the quantization problem of the very few cherry parameters in parameter heterogeneity.
  3. 3.
    How to find the optimal mixed-parameter selection strategy? Based on heterogeneity that distinguishes normal parameters from cherry parameters, the impact-based metric demonstrates the highest discriminative capability.
  4. 4.
    How to quantize parameters according to parameter heterogeneity? We propose CherryQ based on these findings. Extensive empirical results verify its effectiveness in 2-, 3-, 4-bit quantization.
@@ 2 Related Work
Quantization Strategies for LLMs Various quantization strategies have been proposed in the literature to reduce the precision of weights and activations while maintaining acceptable accuracy. These strategies can be broadly classified into post-training quantization and quantization-aware training [14]. Post-training quantization methods, such as OBD, OBS, and GPTQ, directly quantize the pre-trained model without fine-tuning [15, 10, 9]. On the other hand, quantization-aware training methods, such as LLM-QAT [18], incorporate quantization operations into the training process to jointly optimize the quantized model. Some works also explore mixed-precision quantization [12] and adaptive quantization bins [7] to achieve a better trade-off between accuracy and efficiency.
Outliers in Language Model Quantization The idea of modeling parameter outliers in LLM quantization is not new. Exploring outliers mainly includes the perspectives of magnitude [18, 7] and activations [4, 6]. For example, from the magnitude perspective, QLoRA assumes that parameters follow a Gaussian distribution [7] and designs information-theoretically optimal quantized bins based on this assumption. [18] keeps outlier parameters in 16-bit precision. From the activation perspective, [17] migrates the outlier amplifier to subsequent modules through an equivalent transformation. Additionally, SqueezeLLM also measures outliers from the perspective of parameter impact [12]. To the best of our knowledge, our work is the first to systematically reveal the outliers (heterogeneity) of parameter impact across different models, and we show a more pronounced imbalance in parameter impacts compared to magnitudes (§ 6.5). Furthermore, we propose a method to unify outlier (cherry) parameter optimization and normal parameter optimization, addressing the optimization challenges of heterogeneous parameters.
@@ 3 Quantifying the Impact of Parameters on Model Performance
$\mathbf{L}(w_{i}+\Delta)-\mathbf{L}(w_{i})=g_{i}\Delta+\frac{1}{2}\mathbf{H}_{ii}\Delta^{2}+O(\Delta^{2})$ (1)
@@
4 End-to-End Mixed-Precision Quantization
The insights gained from Figure 1 highlight the heterogeneity in model parameters. To mitigate the impact of cherry parameters on quantization, we propose to preserve their high-precision values during the quantization process. By maintaining the fidelity of these critical parameters, we ensure that the essential information they capture is not compromised.
Optimizing mixed-precision parameters in LLMs presents a unique challenge in the widely adopted Post-Training Quantization (PTQ) framework [14]. If we do not allow the updates of the cherry parameters, the quantization will certainly lose the flexibility provided by these critical parameters. This prevents the cherry parameters from reaching their optimum. On the other hand, PTQ struggles to simultaneously optimize high-precision cherry parameters and low-precision normal parameters. This is because the cherry parameter updates during the PTQ process significantly affect the optimal values of the normal parameters. So normal parameters need to be continually updated as the cherry parameter varies. However, in PTQ, once the normal parameters are quantized, they cannot be further updated. This prevents the early-stage quantized parameters from reaching their optimal values.
To address this challenge, we propose a novel approach that end-to-end optimize the mixed-precision parameters via backpropagation. Our method leverages a quantization-aware training framework. To simultaneously optimize both the cherry parameters and normal parameters, we use two separate backpropagation strategies. The high-precision cherry parameters are updated using standard gradient descent, while the low-precision normal parameters employ the Straight-Through Estimator (STE) trick [3] for low-precision gradient descent. This unified backpropagation enables the end-to-end optimization of both cherry parameters and normal parameters, enhancing the overall optimization effect. We show the quantization in Algorithm 1.
Algorithm 1 CherryQ
@@ 4 End-to-End Mixed-Precision Quantization
3:$\mathbf{C}\leftarrow\{w_{i}\in\mathbf{W}\mid\mathbf{H}_{ii}>\tau\}$triangleright Identify cherry parameters
@@
6:     $L\leftarrow\mathrm{model}(x;\mathbf{C}\cup Quant(\mathbf{N}))$triangleright Compute loss w.r.t. mixed-precision parameters
7:     $\mathbf{C}\leftarrow\mathbf{C}-\eta\frac{\partial L}{\partial\mathbf{C}}$triangleright Standard gradient descent for cherry parameters
8:     $\mathbf{N}\leftarrow\mathbf{N}-\eta\cdot\mathrm{STE}(\frac{\partial L}{\partial\mathbf{N}})$triangleright Gradient approximation by STE for normal parameters
@@
5 Heterogeneity-based Cherry Parameter Selection
Correctly identifying cherry parameters is one of the main challenges of CherryQ quantization. Candidate metrics for parameter influences include weights [13, 18, 6], activations [17, 24], and impacts ($\mathbf{H}_{ii}$). We propose that an effective metric should reflect heterogeneity, specifically by differentiating the influence of cherry parameters and normal parameters of the model.
To this end, we define the heterogeneity score. In Figure 1, a small subset of parameters exhibit significantly higher impacts compared to the maximum of the majority. Inspired by this, the heterogeneity score is defined as the ratio of the mean impact of the top 1% parameters to the maximum impact of the bottom 99% parameters, as shown in Equation (4). A higher heterogeneity score indicates a more significant disparity in parameter importance.
Heterogeneity $\text{Heterogeneity Score}(f)=\frac{\text{Mean}(f(w_{i})_{\text{ top 1\%}})}{\text{Max}(f(w_{i})_{\text{ bottom 99\%}})}$ (4)
where $f(w_{i})$ denotes the parameter influence for parameter $w_{i}$, and $f$ is chosen from parameter weights, activations, and impacts.
Figure 2 presents the heterogeneity scores for different metrics across various LLMs. The impact-based metric consistently shows higher heterogeneity scores compared to weights and activations. This indicates that the impact metric better distinguishes between the normal and cherry parameters, thus providing a more effective means of identifying cherry parameters. The validity of using heterogeneity scores for cherry parameter selection will be further verified in Sec 6.5, demonstrating that higher heterogeneity scores lead to better model performance.
@@ 5 Heterogeneity-based Cherry Parameter Selection
Figure 2: Scatter distribution of heterogeneity scores for different parameter matrices in LLMs. Each point represents a parameter matrix.
Data Independence We investigate whether impact-based parameter heterogeneity exhibits data independence - that is, whether different data samples share the same cherry parameters. Applying identical cherry parameters across different samples is only valid when there is data independence. To examine data independence within the same dataset, we randomly selected five sets of 128 WikiText-2 samples. We evaluated the overlap of their cherry parameters in each pair of sample sets. To evaluate cross-dataset data independence, we performed 9 independent sampling trials, each collecting 128 samples from C4 and 128 samples from WikiText-2 to evaluate the overlap. For Vicuna, we further added Sharegpt as a data source. Table 1 presents the overlap ratios of the cherry parameters. Despite the fact that cherry parameters constitute only $1/256$ of the total parameters, all models demonstrate significant overlap ratios. This finding suggests that the cherry parameters possess an inherent data-independent nature.
Table 1: Overlap ratio (%) of cherry parameters (top 1/256), where L2: LLaMA-2, V: Vicuna-1.5, M: Mistral, G: Gemma.
Model L2-7B L2-13B V-7B V-13B M-7B G-7B
Within dataset 84 75 68 63 89 90
Across datasets 68 66 65 60 85 86
@@ 6.1 Implementation Details
Parameter Representation: Based on the observation that cherry parameters occupy a very small proportion, for each row of parameters in each parameter matrix, we consider only the top 1/256 parameters with the highest impact as cherry parameters and retain their FP16 precisions. For example, the parameter matrix size of LLaMA2-7B is $4096\times 4096$. So we average the impact across all rows for each column and then select the top $16$ columns with the highest average impact, resulting in $16\times 4096$ parameters as cherry parameters. Furthermore, to recover the complete parameter matrix, an INT16 is required to record the indices of these $16$ columns. Thus, the storage overhead for the column indices is minimal. For normal parameters, we employ full range symmetric MinMax quantization to quantize their weights. And we adopt a widely-used parameter grouping strategy. For more details, see Sec C.
Quantization Datasets: For the quantization of the base LLMs, we follow [9] to use C4 [20] as the training data. We selected the first four partitions of C4 and chose data with a length of $\geq 2048$ tokens, resulting in a total of 50k samples of 2048 tokens. For the chat LLMs, since Vicuna-1.5 [5] is obtained by supervised fine-tuning based on ShareGPT [5], we also use the ShareGPT dataset for training. We used a total of 20k training samples from ShareGPT for QAT and Cherry.
Baselines We compare our method with various quantization methods, including QAT [18], GPTQ [9], SqueezeLLM [12], OmniQuant [21], and AWQ [17]. For OmniQuant and AWQ, we use their results reported in [21]. For SqueezeLLM, we use the results in its original paper [12]. For GPTQ, its 4-bit model is obtained from the open-source  [^1] https://huggingface.co/TheBloke. Due to the lack of a 3-bit GPTQ model, we quantize the model ourselves via the implementation of Auto-GPTQ  [^2] https://github.com/AutoGPTQ/AutoGPTQ. Since CherryQ is based on QAT, for fair comparisons, the implementation of QAT is the same as CherryQ, except that it does not handle cherry parameters.